OpenStudy (anonymous):

express 4x + 15= 13 in point-slope form

OpenStudy (anonymous):

y=mx+b form solve for y? Where is teh y?

OpenStudy (anonymous):

ug,,, 4x+5y =13

OpenStudy (anonymous):

okay slope for y

OpenStudy (anonymous):

subtract sorry

OpenStudy (anonymous):

4x+5y=13 minus 4x from both sides 5y=-4x+13 divide both sides by 5 \[y=\frac{ -4 }{ 5 }x +\frac{ 13 }{ 5 }\]

OpenStudy (anonymous):

its a hard one

OpenStudy (anonymous):

ohhhhhhhh this is starting to make sense now! (i think)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

usually you don't end with fractions, textbooks give nice easy ones

OpenStudy (anonymous):

not this one,,, but ty!

OpenStudy (anonymous):

nP